OpenStudy (ktbug3):

Combine like terms. 7 + 4x – 10x + 15 A) 22 + 6x B) 22 – 14x C) 22 – 6x D) 11x – 5x will fan and medal

jabez177 (jabez177):

4 - 10 = ?

OpenStudy (anonymous):

Do you know how to solve for this? I can show you if you would like!

OpenStudy (ktbug3):

-6

OpenStudy (ktbug3):

@itsbribro thankyou pls do:)

OpenStudy (anonymous):

Here are the steps: 7+4x−10x+15 =7+4x+−10x+15 Now lets combine the like terms of this equation. =7+4x+−10x+15 =(4x+−10x)+(7+15) and now, from these steps, what do you think the answer is?

OpenStudy (gabbyalicorn):

Simplify 7+4x−10x+15 =7+4x+−10x+15 Combine Like Terms =7+4x+−10x+15 =(4x+−10x)+(7+15) =−6x+22 Answer: =−6x+22 A ^

OpenStudy (anonymous):

−6x+22 is your answer. Gosh Gabby :P

OpenStudy (anonymous):

Gotta steal all my questions >.< XD Thanks! :P I will give you a medal! :D

OpenStudy (gabbyalicorn):

Sorry, didn't mean to steal your thunder. ^///^ I like answering algebraic equations.

OpenStudy (anonymous):

Same! :D